Quantitative Aptitude Practice Question and Answer
8Q: In an old Palace there are 50 stairs to go on top floor. Each is 20 cm high, 30 cm wide and 60 cm long. If all the stairs are re-plastered at the rate of Rs. 4 per 100 sq.cm. What is the expenditure ? 1972 05b5cc6cbe4d2b4197774e0f9
5b5cc6cbe4d2b4197774e0f9- 1Rs. 4680false
- 2Rs. 7600false
- 3Rs. 6000true
- 4Rs. 5480false
- Show AnswerHide Answer
- Workspace
- SingleChoice
Answer : 3. "Rs. 6000"
Explanation :
Answer: C) Rs. 6000 Explanation: Total are of all stairs = 50 (20x60 + 30x60) = 50 (1200 + 1800) = 50 x 3000 = 150000 sq cm. Given rate of 100 sq cm = Rs. 4 Then for 150000 sq. cm = ? => 4 x 150000/100 = Rs. 6000
Q: A number, when 35 is subtracted from it, reduces to its 80 percent. What is four - fifth of that number ? 1969 05b5cc6dee4d2b4197774ea4b
5b5cc6dee4d2b4197774ea4b- 1175false
- 2140true
- 3135false
- 495false
- Show AnswerHide Answer
- Workspace
- SingleChoice
Answer : 2. "140"
Explanation :
Answer: B) 140 Explanation: x - 35 = 80x/100 => x = 175 => 4x/5 = 4x175/5 = 140.
Q: In a box, there are 9 blue, 6 white and some black stones. A stone is randomly selected and the probability that the stone is black is ¼. Find the total number of stones in the box? 1969 05b5cc64ae4d2b4197774bc47
5b5cc64ae4d2b4197774bc47- 115false
- 218false
- 320true
- 424false
- Show AnswerHide Answer
- Workspace
- SingleChoice
Answer : 3. "20"
Explanation :
Answer: C) 20 Explanation: We know that, Total probability = 1 Given probability of black stones = 1/4 => Probability of blue and white stones = 1 - 1/4 = 3/4 But, given blue + white stones = 9 + 6 = 15 Hence, 3/4 ----- 15 1 ----- ? => 15 x 4/3 = 20. Hence, total number of stones in the box = 20.
Q: A is two years older than B who is twice as old as C. If the total of the ages of A, B and C be 27, then how old is B ? 1969 05b5cc6e7e4d2b4197774ed6f
5b5cc6e7e4d2b4197774ed6f- 110 yrstrue
- 211 yrsfalse
- 312 yrsfalse
- 413 yrsfalse
- Show AnswerHide Answer
- Workspace
- SingleChoice
Answer : 1. "10 yrs"
Explanation :
Answer: A) 10 yrs Explanation: Let C's age be x years. Then, B's age = 2x years. A's age = (2x + 2) years. (2x + 2) + 2x + x = 27 5x = 25 => x = 5 Hence, B's age = 2x = 10 years.
Q: In a triangle ABC, angle 'a' is the smallest one & A+7B = 112. Find the range of value 'c' ? 1969 05b5cc6f4e4d2b4197774efe6
5b5cc6f4e4d2b4197774efe6- 1138 to 164false
- 2125 to 147false
- 3148 to 158false
- 4152 to 164true
- Show AnswerHide Answer
- Workspace
- SingleChoice
Answer : 4. "152 to 164"
Explanation :
Answer: D) 152 to 164 Explanation: A+7B=112 it is clear that A = 14, then it becomes A>=B, but A is smallest angle) givenso, range of A is 0.0001 to 13.9999 ( I am taking upto 4 decimal places)so, range of B becomes 14 to 16 ( after rounding off to 4 decimal places)so, range of C becomes 152 to 164 ( after rounding off to 4 decimal places)
Q: The total of the ages of Ashok, Akbar and Anil is 65 years. What was the total of their ages six years ago ? 1968 05b5cc723e4d2b4197774f695
5b5cc723e4d2b4197774f695- 159false
- 243false
- 357false
- 447true
- Show AnswerHide Answer
- Workspace
- SingleChoice
Answer : 4. "47"
Explanation :
Answer: D) 47 Explanation: Let their ages be x, y, zHence given that x + y + z = 65Now, six years ago x age is x - 6, y is y - 6 and that of z is z - 6Then the sum of their ages six years ago is (x + y + z) - (6 + 6 + 6) = 65 - 18 = 47
Q: Mohan spends 23% of an amount of money on an insurance policy, 33% on food, 19% on children’s education and 16% on recreation. He deposits the remaining amount of Rs. 504 in bank. How much amount does he spend on food and insurance policy together? 1967 05b5cc678e4d2b4197774c160
5b5cc678e4d2b4197774c160- 1Rs. 2326false
- 2Rs. 5876false
- 3Rs. 1741false
- 4Rs. 3136true
- Show AnswerHide Answer
- Workspace
- SingleChoice
Answer : 4. "Rs. 3136"
Explanation :
Answer: D) Rs. 3136 Explanation: Let total amount = Rs. P Total expenditure = (23 + 33 + 19 + 16)% of P = 91%ofP Remaining money = (100 - 91)% of P = 9% of P But given in the question that, 9% of P = 504 9p/100 = 504 => P = 50400/9 => P = Rs. 5600 Required money spent on Food & Insurance = 23 + 33 % = 56% of 5600 = 56 x 5600/100 = 56 x 56 = Rs. 3136
Q: A card is drawn from a pack of 52 cards. The probability of getting a queen of the club or a king of the heart is? 1967 05b5cc6bde4d2b4197774d94c
5b5cc6bde4d2b4197774d94c- 11/26true
- 21/13false
- 32/13false
- 41/52false
- Show AnswerHide Answer
- Workspace
- SingleChoice
Answer : 1. "1/26"
Explanation :
Answer: A) 1/26 Explanation: Here in this pack of cards, n(S) = 52 Let E = event of getting a queen of the club or a king of the heart Then, n(E) = 2 P(E) = n(E)/n(S) = 2/52 = 1/26

